OpenStudy (anonymous):

The number A of tablet complrters sold (in millions) can be modeled by the functionA : 4.5t2 + 43.5t * 17, where /represents the year after 2010. A . In what year did the tablet computer sales reach 65 million? B . F'ind the average rate of change from 2010 to 2012 and interpret the meaning in the context of the situation. C . Do you think this model will be accurate after a new, innovative computer is developed? Explain.

OpenStudy (freckles):

So you want to know when A=65 solve the equation 65=4.5t^2+43.5t*17 (if that is the right equation; kind of wonder if you meant to put +17 at then end instead of *17)

OpenStudy (freckles):

for B evaluate: \[\frac{f(2)-f(0)}{2-0} \\ \text{ since } t=2 \text{ actually means 2 years after 2010 which is the year 2012}\\ \text{ and since } t=0 \text{ means the year 2010}\]

OpenStudy (anonymous):

I substitute a = 65 I solve t=1 Then 1+ 2010 = 2011 ?

OpenStudy (freckles):

right I don't think they care about what happened before 2010... you know because you do get also t=-10.6667 as a solution

OpenStudy (freckles):

which I would take that to mean like little more 10 years before but like I said I don't think they care about what happens before

OpenStudy (anonymous):

how I do B?

OpenStudy (freckles):

oops those f were suppose to be A's

OpenStudy (freckles):

sorry evaluate: \[\frac{A(2)-A(0)}{2-0}\] there does that make more sense ?

OpenStudy (freckles):

That is the average rate of change of A from the year 2010 to 2012 (aka t=0 to t=2)

OpenStudy (anonymous):

ok I see

OpenStudy (anonymous):

hanks

OpenStudy (freckles):

np

OpenStudy (anonymous):

C?

OpenStudy (freckles):

C) I don't know...The economy is always changing...But I think they are making computers cheaper... There will be more people. Some people will be fine with their old computer and won't think it necessary to purchase another computer.

OpenStudy (anonymous):

Thanks

OpenStudy (freckles):

That question might require some research.

OpenStudy (freckles):

For example I don't even know if the function they gave is a great model currently for any computer type available without doing any research.

OpenStudy (anonymous):

ok, I will try search for C, Thanks

OpenStudy (freckles):

what do you think @dan815

OpenStudy (anonymous):

I no idea I have to go , thanks bye

OpenStudy (dan815):

sry i wasnt here